4 of 16> Solve for x. 375 5x3 Express the answer to the hundredths place (i.e., two digits after the decimal point). ho 80 144
Answer #1

375 = 5 x^3

x^3 = 375/5 = 75

x = (75)^(1/3)

Now we have to find cuberoot of 75.

(4^3 = 64

5^3 = 125)


x = 4.22 ........ans